# Neural networks for the approximation of Euler’s elastica

### v1

@@ 4 The discrete network
The discretisation of Euler’s elastica presented in Section 2.1 provides discrete solutions on a set of nodes along the curve. These solutions can sometimes be hard to obtain since a non-convex optimisation problem needs to be solved, and the number of nodes can be large. This motivates the use of neural networks to learn the approximate solution on the internal nodes, for a given set of boundary conditions. The data set $\Omega$ consists of $M$ precomputed discrete solutions
@@
$\mathbf{x}^{i}=\left(\mathbf{q}^{i}_{0},\mathbf{q}^{i}_{N},(\mathbf{q}^{i}_{0})^{\prime},(\mathbf{q}^{i}_{N})^{\prime}\right)\in\mathbb{R}^{8}$
@@
$\mathbf{y}^{i}=(\hat{\mathbf{q}}_{1}^{i},\ldots,\hat{\mathbf{q}}_{N-1}^{i},(\hat{\mathbf{q}}^{i}_{1})^{\prime},\ldots,(\hat{\mathbf{q}}^{i}_{N-1})^{\prime}))\in\mathbb{R}^{4(N-1)}$
are the computed solutions at the internal nodes that serve as output data for the training of the network.
For any symmetric positive definite matrix $W$, we define the weighted norm $\|\mathbf{x}\|_{W}=\|W\mathbf{x}\|_{2}$. The weighted MSE loss
$\mathrm{Loss}(\boldsymbol{\rho})=\frac{1}{4M(N-1)}\sum_{i=1}^{M}\left\|q_{\boldsymbol{\rho}}^{\textrm{d}}\left(\mathbf{x}^{i}\right)-\mathbf{y}^{i}\right\|_{W}^{2},$ (11)
will be used to learn the input-to output map $q_{\boldsymbol{\rho}}^{\textrm{d}}:\mathbb{R}^{8}\to\mathbb{R}^{4(N-1)}$, where the superscript $\mathrm{d}$ stands for discrete. One should be aware that there is a numerical error in $\mathbf{y}^{i}$ compared to the exact solution and the size of this error will pose a limit to the accuracy of the neural network approximation.
@@ 4.1 Numerical experiments
This section provides experimental support to the proposed learning framework. We perform a series of experiments varying the architecture of the neural network and the hyperparameters in the training procedure. The codes to run the experiments in this work are written using the machine learning library PyTorch [45]. We use the Adam optimiser [44] for the training, carefully selecting learning rate and weight decay to prevent over-fitting, see Table 2. In (11) we use the weight matrix
$W=I+\gamma G^{T}G$
where $G=S-I$ with $S$ the forward shift operator on vectors of $\mathbb{R}^{4(N-1)}$. We test a range of different batch sizes $B\leq M$ and fix the total number of epochs to $300$. Finally, we also test for the influence of performing normalisation of the input data. We collect in Table 2 all the hyperparameters and network architectures with their corresponding ranges.
Hyperparameter
Range
Distribution
architecture
{MLP, ResNet}
discrete uniform
normalisation
{True, False}
discrete uniform
activation function $\sigma$
{Tanh, Swish, Sigmoid, ReLU, LeakyReLU}
discrete uniform
#layers $\ell$
$\{0,1,2,3,4\}$
discrete uniform
#hidden nodes in each layer
$[32,1024]\cap\mathbb{N}$
discrete uniform
learning rate $\eta$
$[1\cdot 10^{-4},1\cdot 10^{-1}]$
log uniform
weight decay
$[1\cdot 10^{-7},5\cdot 10^{-4}]$
log uniform
$\gamma$
$[0,1\cdot 10^{-2}]$
uniform
batch size
$\{32,64,128\}$
discrete uniform
Table 2: Hyperparameter ranges for the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends data set. The first column of the table reports the hyperparameters and network architectures we test for. The second describes the set of allowed values for each, while the third specifies how such values are explored through Optuna. MLP corresponds to the network in Equation (9), Section 3, while ResNet is a residual neural network defined in Equation (19) of Appendix A.
@@
We rely on the software framework Optuna [46] to automate and efficiently conduct the search for the combination that yields the best result. This is reported in Table 3. The resulting training error on the both-end data set is $2.791\cdot 10^{-7}$, and the test error on a set of trajectories belonging to the same data set is $3.028\cdot 10^{-7}$. Figure 1 compares test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$. We remark that, as already clear from the low value of the training and test errors, the network can accurately replicate the behaviour of the training and test data. Furthermore, since the network is trained only on the internal nodes and the boundary values are appended to the predicted solution in a post processing phase, we have zero errors at the end nodes. On the other hand, since this discrete approach does not relate the components as evaluations of a smooth curve, there is no regular behaviour in the error.
Selected hyperparameters
architecture
MLP
normalisation
True
activation function $\sigma$
Tanh
#layers $\ell$
4
#hidden nodes in each layer
879
learning rate $\eta$
$1.378\cdot 10^{-3}$
weight decay
$1.535\cdot 10^{-7}$
$\gamma$
$4.242\cdot 10^{-3}$
batch size
32
Table 3: Combination of hyperparameters yielding the best results for the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends data set with $90\%\--10\%$ splitting into training and test set. The results are shown in Figure 1.
@@
Figure 1: Comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$ for the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends data set with $90\%\--10\%$ splitting into training and test set. These results are obtained with the hyperparameters from Table 3, that yield a training error equal to $2.791\cdot 10^{-7}$ and a test error equal to $3.028\cdot 10^{-7}$ For presentation purposes, only 10 randomly selected trajectories are considered in the first two plots.
@@
As an additional evaluation of the deep learning framework’s behaviour, we conduct experiments to assess how the learning process performs when the number of training data varies, i.e., with different splittings of the data set into training and test sets. We report the results in Table 4 and summarise the corresponding hyperparameters in Table 12 of the Appendix.
@@
Data set splittingTraining - $\begin{array}[]{c}\text{Data set splitting}\\ \text{Training \-- test}\end{array}$
Training accuracy Test accuracy
10% - 10%
$6.530\cdot 10^{-5}$
$5.636\cdot 10^{-4}$
20% - 10%
$2.096\cdot 10^{-5}$
$3.457\cdot 10^{-5}$
40% - 10%
$2.186\cdot 10^{-6}$
$3.494\cdot 10^{-6}$
90% - 10%
$2.791\cdot 10^{-7}$
$3.028\cdot 10^{-7}$
@@
Table 4: Behaviour of the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends data set with fewer training data points. The size of the training set varies, while that of the test set is fixed. The last row corresponds to the results in Figure 1.
We also report results obtained by merging the both-end and the right-end trajectories, with $90\%\--10\%$ splitting of the whole new data set into training and test set. The results are shown in Figure 2 and the selected hyperparameters are collected in Table 5. The resulting training and test errors are, respectively, $3.047\cdot 10^{-7}$ and $3.141\cdot 10^{-7}$. Finally, we remark that the test accuracy is a good measure of the generalisation error of neural network under the hypothesis that the test and the training sets are independent of each other, but follow the same distribution. If we test over input boundary conditions that not only are unseen during the training, but also do not belong in the the same range of the training data, the resulting accuracy is expected to be low, since neural networks are in general not able to perform this sort of extrapolation. To show this, we consider the neural network trained and tested over the both-ends data set, related to the results in Figure 1 and in the last row of Table 4. We use 10% of the right-end data set as a test set, and we obtain a test error equal to $2.228\cdot 10^{-2}$. This highlights that care must be taken when using the trained network to make inference over new input data.
@@
Figure 2: Comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$ for the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends + right-end data set with $90\%\--10\%$ splitting into training and test set. These results are obtained with the hyperparameters from Table 5, that yield a training error equal to $3.047\cdot 10^{-7}$ and a test error equal to $3.141\cdot 10^{-7}$. For presentation purposes, only 10 randomly selected trajectories are considered in the first two plots.
Selected hyperparameters
architecture
MLP
normalisation
True
activation function $\sigma$
LeakyReLU
#layers $\ell$
2
#hidden nodes in each layer
1006
learning rate $\eta$
$3.611\cdot 10^{-3}$
weight decay
$1.515\cdot 10^{-7}$
$\gamma$
$6.388\cdot 10^{-3}$
batch size
32
Table 5: Combination of hyperparameters yielding the best results for the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends + right-end data set with $90\%\--10\%$ splitting into training and test set. The results are shown in Figure 2.
@@ 5 The continuous network
The approach described in the previous section shows accurate results, given a large enough amount of beam discretisations with a fixed number of nodes $N+1$, equally distributed in $[0,L]$. It seems reasonable to expect that the parametric model’s approximation quality improves when the number of discretisation nodes increases. However, in this approach, the dimension of the predicted vector grows with $N$, and hence minimising the loss function (11) becomes more difficult. In addition, the fact that the discrete network approach depends on the spatial discretisation of the training data restricts the output dimension to a specific number of nodes. Consequently, there would be two main options to assess the solution at different locations: training the network once more, or interpolating the previously obtained approximation. These limitations make such a discrete approach less appealing and suggest that having a neural network that is a smooth function of the arc length coordinate $s$ can be beneficial. This modelling assumption would also be compatible with different discretisations of the curve and would not suffer from the curse of dimensionality if more nodes were added. In this setting, the discrete node $s_{k}$ at which an approximation of the solution is available, is included in the input data together with the boundary conditions. As a result, we work with the following data set
@@
$\mathbf{x}^{i}=\left(\,\mathbf{q}^{i}_{0},\,\mathbf{q}^{i}_{N},\,(\mathbf{q}^{i}_{0})^{\prime},\,(\mathbf{q}^{i}_{N})^{\prime}\right)\in\mathbb{R}^{8},$
@@
$\mathcal{D}:\mathcal{C}^{\infty}\left([0,L],\mathbb{R}^{2}\right)\to\mathcal{C}^{\infty}\left([0,L],\mathbb{R}^{2}\right),\,\,\mathcal{D}\left(q_{\boldsymbol{\rho}}^{\mathrm{c}}\left(\mathbf{x}^{i}\right)\right)(s_{k})=\frac{d}{ds}\left(q_{\boldsymbol{\rho}}^{\rm c}(\boldsymbol{x}^{i})\right)(s)\Big{|}_{s=s_{k}},$
@@
To train the network $q_{\boldsymbol{\rho}}^{c}$, we define the loss function
$\textrm{Loss}(\boldsymbol{\rho})=\frac{1}{4M(N+1)}\sum_{i=1}^{M}\sum_{k=0}^{N}\left(\left\|{y_{\boldsymbol{\rho}}}\left(\mathbf{x}^{i}\right)(s_{k})-{\boldsymbol{y}}_{k}^{i}\right\|^{2}_{2}+\gamma\left(\left\|\pi_{\mathcal{D}}\left(y_{\boldsymbol{\rho}}\left(\mathbf{x}^{i}\right)(s_{k})\right)\right\|_{2}^{2}-1\right)^{2}\right),$ (12)
where $\pi_{\mathcal{D}}:\mathbb{R}^{8}\to\mathbb{R}^{4}$ is the projection on the second component $\mathcal{D}(q_{\boldsymbol{\rho}}^{\rm c}(\boldsymbol{x}^{i}))(s_{k})$, and $\gamma\geq 0$ weighs the violation of the normality constraint. The map $q_{\boldsymbol{\rho}}^{\textrm{c}}$ is now a neural network that associates each set of boundary conditions $\boldsymbol{x}^{i}$ with a smooth curve $q_{\boldsymbol{\rho}}^{\textrm{c}}\left(\boldsymbol{x}^{i}\right):[0,L]\to\mathbb{R}^{2}$ that can be evaluated at every point $s\in[0,L]$. We denote this network with the superscript $\mathrm{c}$, since this curve is in particular continuous. The outputs $q_{\boldsymbol{\rho}}^{\textrm{c}}\left(\boldsymbol{x}^{i}\right)(s)\in\mathbb{R}^{2}$ are approximations of the configuration of the beam at $s\in[0,L]$.
We point out that, contrarily to the discrete case, here we learn approximations of $\mathbf{q}(s)$ also on the end nodes, i.e., at $s=0$ and $s=L$. This is due to the fact that we do not impose the boundary conditions by construction. Even though there are multiple approaches to embed them into the network architecture, the one we try in our experiments made the optimisation problem too difficult, thus we only impose the boundary conditions weakly in the loss function.
@@
is a neural network and the function $\pi:\mathbb{R}^{8}\to\mathbb{R}^{2}$ extracts the tangential angles from the boundary conditions, i.e., $\pi\left(\boldsymbol{x}^{i}\right)=\left(\theta_{0}^{i},\theta_{N}^{i}\right)$. Such a network should approximate the angular function $\theta:[0,L]\ni s\to\mathbb{R}$, so that
$\tau_{\boldsymbol{\rho}}^{\mathrm{c}}\left(\boldsymbol{x}^{i}\right)(s):=\left(\cos\left(\theta_{\boldsymbol{\rho}}^{\textrm{c}}\left(\boldsymbol{x}^{i}\right)(s)\right),\sin\left(\theta_{\boldsymbol{\rho}}^{\textrm{c}}\left(\boldsymbol{x}^{i}\right)(s)\right)\right)\in\mathbb{R}^{2}$ (14)
@@
$q_{\boldsymbol{\rho}}^{\textrm{c}}\left(\boldsymbol{x}^{i}\right)(s)=\mathbf{q}_{0}+\mathcal{I}\left(\tau_{\rho}^{\mathrm{c}}\left(\boldsymbol{x}^{i}\right)\right)(s),$ (15)
@@
$\mathcal{I}\left(\tau_{\rho}^{\mathrm{c}}\left(\boldsymbol{x}^{i}\right)\right)(s)\approx\int_{0}^{s}\tau_{\rho}^{\mathrm{c}}\left(\mathbf{x}^{i}\right)(\bar{s})\mathrm{d}\bar{s}.$
@@
$y_{\boldsymbol{\rho}}\left(\mathbf{x}^{i}\right)(s_{k}):=\left(q_{\boldsymbol{\rho}}^{\mathrm{c}}\left(\mathbf{x}^{i}\right)(s_{k}),\,\tau_{\boldsymbol{\rho}}^{\rm c}\left(\boldsymbol{x}^{i}\right)(s_{k})\right),$ (16)
with components defined as in Equations (14) and (15). This allows us to train the network $\theta_{\boldsymbol{\rho}}^{\rm c}$ by minimising the same loss function as in Equation (12), where this time $y_{\boldsymbol{\rho}}^{\rm c}$ is given by Equation (16). Furthermore, since by construction this case satisfies $\left\|\pi_{\mathcal{D}}\left(y_{\boldsymbol{\rho}}^{\rm c}(\boldsymbol{x}^{i})(s)\right)\right\|_{2}=\left\|\tau_{\boldsymbol{\rho}}^{\mathrm{c}}\left(\boldsymbol{x}^{i}\right)(s)\right\|_{2}\equiv 1$, we set $\gamma=0$. We present numerical experiments for the two proposed continuous networks $q_{\boldsymbol{\rho}}^{\textrm{c}}$ and $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$. In the latter case, by neural network architecture we refer to $\hat{\theta}_{\boldsymbol{\rho}}^{\rm c}$ rather than $\theta_{\boldsymbol{\rho}}^{\rm c}$ in what follows. We analyse $q_{\boldsymbol{\rho}}^{\textrm{c}}$ more thoroughly in Section 5.1, mirroring most of the discrete case experiments. In Section 5.2 we study how the results are affected when we impose the arc length parametrization and enforce the boundary conditions to be exactly satisfied by the network $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$.

### v2

@@ 4 The discrete network
The discretisation of Euler’s elastica presented in Section 2.1 provides discrete solutions on a set of nodes along the curve. These solutions can sometimes be hard to obtain since a global optimisation problem needs to be solved, and the number of nodes can be large. This motivates using neural networks to learn the approximate solution on the internal nodes for a given set of boundary conditions. The data set $\Omega$ consists of $M$ precomputed discrete solutions
@@
$\mathbf{x}^{i}=\left(\mathbf{q}^{i}_{0},(\mathbf{q}^{i}_{0})^{\prime},\mathbf{q}^{i}_{N},(\mathbf{q}^{i}_{N})^{\prime}\right)\in\mathbb{R}^{8}$
@@
$\mathbf{y}^{i}=(\hat{\mathbf{q}}_{1}^{i},(\hat{\mathbf{q}}^{i}_{1})^{\prime},\ldots,\hat{\mathbf{q}}_{N-1}^{i},(\hat{\mathbf{q}}^{i}_{N-1})^{\prime})\in\mathbb{R}^{4(N-1)}$
are the computed solutions at the internal nodes that serve as output data for the network’s training.
For any symmetric positive definite matrix $W$, we define the weighted norm $\|\mathbf{x}\|_{W}^{2}=\mathbf{x}^{\top}W\mathbf{x}$. The weighted MSE loss
$\mathrm{Loss}(\boldsymbol{\rho})=\frac{1}{4M(N-1)}\sum_{i=1}^{M}\left\|q_{\boldsymbol{\rho}}^{\textrm{d}}\left(\mathbf{x}^{i}\right)-\mathbf{y}^{i}\right\|_{W}^{2}$ (11)
will be used to learn the input-to-output map $q_{\boldsymbol{\rho}}^{\textrm{d}}:\mathbb{R}^{8}\to\mathbb{R}^{4(N-1)}$, where the superscript $\mathrm{d}$ stands for discrete. One should be aware that there is a numerical error in $\mathbf{y}^{i}$ compared to the exact solution and the size of this error will pose a limit to the accuracy of the neural network approximation.
@@ 4.1 Numerical experiments
This section provides experimental support to the proposed learning framework using the machine learning library PyTorch [49]. The experiments of this section are run on a CPU machine. We perform a series of experiments varying some hyperparameters in the training procedure. We fix the batch size $B$ to 32 and use the Adam optimiser [48] for the training with learning rate $10^{-3}$ and weight decay set to 00. In (11) we use the weight matrix
$W=I+\gamma G^{\top}G,$
@@
where $G=S^{4}-I$ with $S$ the forward shift operator on vectors of $\mathbb{R}^{4(N-1)}$. This choice of $G$ allows us to compute differences between corresponding entries of the input associated with neighbouring nodes. We determine the number of epochs for training both the discrete and continuous networks based on experimental evidence. We fix a high enough number which allows us to achieve qualitatively accurate predictions and ensure that both training and validation losses start to plateau a few epochs before the set maximum. We consider a multi-layer perceptron with the hyperbolic tangent as an activation function, and we vary the number of layers and the number of hidden nodes in each layer. We also test different values of the parameter $\gamma$ in the weight matrix $W$. We rely on the software framework Optuna [50], which employs Bayesian optimisation methods to automate and efficiently conduct the search for the combination that yields the best result. We collect in Table 4 the hyperparameters with the corresponding ranges and in Table 5 the selected values. The resulting training error on the both-end data set is $1.14\cdot 10^{-7}$, the validation error is $2.151\cdot 10^{-7}$, and the test error is $4.009\cdot 10^{-7}$. Figure 1 compares test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$. We remark that, as already clear from the low value of the training and test errors, the network can accurately replicate the behaviour of the training and test data. Furthermore, we have zero errors at the end nodes since the network is trained only on the internal nodes and the boundary values are appended to the predicted solution in a post-processing phase. On the other hand, since this discrete approach does not relate the components as evaluations of a smooth curve, there is no regular behaviour in the error.
@@
Figure 1: Comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$ for the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends data set with $80\%\--10\%\--10\%$ splitting into training, validation, and test sets. The mean squared error on the test set equals $4.009\cdot 10^{-7}$. For presentation purposes, only 10 randomly selected trajectories are considered in the first two plots.
@@
As an additional evaluation of the deep learning framework’s behaviour, we conduct experiments to assess how the learning process performs when the number of training data varies, i.e., with different splittings of the data set into training, validation, and test sets. We report the results in Table 2 and summarise the corresponding hyperparameters in Table 5 of the Appendix.
@@
Data set splittingTraining - validation - $\begin{array}[]{c}\text{Data set splitting}\\ \text{Training \-- validation \-- test}\end{array}$ Training accuracy Validation accuracy Test accuracy
10% - 10% - 10% $2.331\cdot 10^{-5}$ $3.874\cdot 10^{-5}$ $8.545\cdot 10^{-4}$
20% - 10% - 10% $1.852\cdot 10^{-6}$ $1.327\cdot 10^{-6}$ $1.361\cdot 10^{-4}$
40% - 10% - 10% $4.802\cdot 10^{-7}$ $4.793\cdot 10^{-7}$ $1.295\cdot 10^{-6}$
80% - 10% - 10% $1.140\cdot 10^{-7}$ $2.151\cdot 10^{-7}$ $4.009\cdot 10^{-7}$
@@
Table 2: Behaviour of the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends data set with fewer training data points. The size of the training set varies, while that of the validation and the test sets is fixed. The last row corresponds to the results in Figure 1.
We also report results obtained by merging the both-end and the right-end trajectories, with $80\%\--10\%\--10\%$ splitting of the whole new data set into training, validation, and test sets. The results are shown in Figure 2 and are obtained with 3 layers, 616 hidden nodes, and $\gamma=7.323\cdot 10^{-3}$. The resulting training, validation, and test errors are, respectively, $9.893\cdot 10^{-8}$, $1.126\cdot 10^{-7}$, and $7.854\cdot 10^{-8}$.
@@
Figure 2: Comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$ for the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends + right-end data set with $80\%\--10\%\--10\%$ splitting into training, validation, and test sets. The mean squared error on the test set equals $7.854\cdot 10^{-8}$. For presentation purposes, only 10 randomly selected trajectories are considered in the first two plots.
@@ 5 The continuous network
The approach described in the previous section shows accurate results, given a large enough amount of beam discretisations with a fixed number of nodes $N+1$, equally distributed in $[0,L]$. It seems reasonable to expect the parametric model’s approximation quality to improve when the number of discretisation nodes increases. However, in this approach, the dimension of the predicted vector grows with $N$, and hence minimising the loss function (11) becomes more difficult. In addition, the fact that the discrete network approach depends on the spatial discretisation of the training data restricts the output dimension to a specific number of nodes. Consequently, there would be two main options to assess the solution at different locations: training the network once more, or interpolating the previously obtained approximation. These limitations make such a discrete approach less appealing and suggest that having a neural network that is a smooth function of the arc length coordinate $s$ can be beneficial. This modelling assumption would also be compatible with different discretisations of the curve and would not suffer from the curse of dimensionality if more nodes were added. In this setting, the discrete node $s_{k}$ at which an approximation of the solution is available, is included in the input data together with the boundary conditions. As a result, we work with the following data set
@@
$\mathbf{x}^{i}=\left(\,\mathbf{q}^{i}_{0},\,(\mathbf{q}^{i}_{0})^{\prime},\,\mathbf{q}^{i}_{N},\,(\mathbf{q}^{i}_{N})^{\prime}\right)\in\mathbb{R}^{8},$
@@
$\mathcal{D}:\mathcal{C}^{\infty}\left([0,L],\mathbb{R}^{2}\right)\to\mathcal{C}^{\infty}\left([0,L],\mathbb{R}^{2}\right),\,\,\mathcal{D}\left(q_{\boldsymbol{\rho}}^{\mathrm{c}}\left(\mathbf{x}^{i}\right)\right)(s_{k})=\frac{d}{ds}\left(q_{\boldsymbol{\rho}}^{\rm c}(\mathbf{x}^{i})\right)(s)\Big{|}_{s=s_{k}},$
@@
To train the network $q_{\boldsymbol{\rho}}^{\mathrm{c}}$, we define the loss function
$\textrm{Loss}(\boldsymbol{\rho})=$ $\frac{1}{4M(N+1)}\sum_{i=1}^{M}\sum_{k=0}^{N}\left(\left\|{y_{\boldsymbol{\rho}}}\left(\mathbf{x}^{i}\right)(s_{k})-{\boldsymbol{y}}_{k}^{i}\right\|^{2}_{2}+\right.$ (12)
$\left.\gamma\left(\left\|\pi_{\mathcal{D}}\left(y_{\boldsymbol{\rho}}\left(\mathbf{x}^{i}\right)(s_{k})\right)\right\|_{2}^{2}-1\right)^{2}\right),$
where $\pi_{\mathcal{D}}:\mathbb{R}^{8}\to\mathbb{R}^{4}$ is the projection on the second component $\mathcal{D}(q_{\boldsymbol{\rho}}^{\rm c}(\mathbf{x}^{i}))(s_{k})$, and $\gamma\geq 0$ weighs the violation of the normality constraint. The map $q_{\boldsymbol{\rho}}^{\textrm{c}}$ is now a neural network that associates each set of boundary conditions $\mathbf{x}^{i}$ with a smooth curve $q_{\boldsymbol{\rho}}^{\textrm{c}}\left(\mathbf{x}^{i}\right):[0,L]\to\mathbb{R}^{2}$ that can be evaluated at every point $s\in[0,L]$. We denote this network with the superscript $\mathrm{c}$ since this curve is, in particular, continuous. The outputs $q_{\boldsymbol{\rho}}^{\textrm{c}}\left(\mathbf{x}^{i}\right)(s)\in\mathbb{R}^{2}$ are approximations of the configuration of the beam at $s\in[0,L]$.
We point out that, contrary to the discrete case, we learn approximations of $\mathbf{q}(s)$ also on the end nodes, i.e., at $s=0$ and $s=L$. This is because we do not impose the boundary conditions by construction. Even though there are multiple approaches to embed them into the network architecture, the one we try in our experiments made the optimisation problem too complex, thus we only impose the boundary conditions weakly in the loss function.
@@
is a neural network, and the function $\pi:\mathbb{R}^{8}\to\mathbb{R}^{2}$ extracts the tangential angles from the boundary conditions, i.e., $\pi\left(\mathbf{x}^{i}\right)=\left(\theta_{0}^{i},\theta_{N}^{i}\right)$. Such a network should approximate the angular function $\theta:[0,L]\ni s\to\mathbb{R}$, so that
$\tau_{\boldsymbol{\rho}}^{\mathrm{c}}\left(\mathbf{x}^{i}\right)(s):=\left(\cos\left(\theta_{\boldsymbol{\rho}}^{\textrm{c}}\left(\mathbf{x}^{i}\right)(s)\right),\sin\left(\theta_{\boldsymbol{\rho}}^{\textrm{c}}\left(\mathbf{x}^{i}\right)(s)\right)\right)\in\mathbb{R}^{2}$ (14)
@@
$q_{\boldsymbol{\rho}}^{\textrm{c}}\left(\mathbf{x}^{i}\right)(s)=\mathbf{q}_{0}+\mathcal{I}\left(\tau_{\rho}^{\mathrm{c}}\left(\mathbf{x}^{i}\right)\right)(s),$ (15)
@@
$\mathcal{I}\left(\tau_{\rho}^{\mathrm{c}}\left(\mathbf{x}^{i}\right)\right)(s)\approx\int_{0}^{s}\tau_{\rho}^{\mathrm{c}}\left(\mathbf{x}^{i}\right)(\bar{s})\mathrm{d}\bar{s}.$
@@
$y_{\boldsymbol{\rho}}\left(\mathbf{x}^{i}\right)(s_{k}):=\left(q_{\boldsymbol{\rho}}^{\mathrm{c}}\left(\mathbf{x}^{i}\right)(s_{k}),\,\tau_{\boldsymbol{\rho}}^{\rm c}\left(\mathbf{x}^{i}\right)(s_{k})\right),$ (16)
with components defined as in Equations (14) and (15). This allows us to train the network $\theta_{\boldsymbol{\rho}}^{\rm c}$ by minimising the same loss function as in Equation (12), where this time $y_{\boldsymbol{\rho}}^{\rm c}$ is given by Equation (16). Furthermore, since by construction this case satisfies $\left\|\pi_{\mathcal{D}}\left(y_{\boldsymbol{\rho}}^{\rm c}(\mathbf{x}^{i})(s)\right)\right\|_{2}=\left\|\tau_{\boldsymbol{\rho}}^{\mathrm{c}}\left(\mathbf{x}^{i}\right)(s)\right\|_{2}\equiv 1$, we set $\gamma=0$. We present numerical experiments for the two proposed continuous networks $q_{\boldsymbol{\rho}}^{\textrm{c}}$ and $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$. In the latter case, by neural network architecture, we refer to $\hat{\theta}_{\boldsymbol{\rho}}^{\rm c}$ rather than $\theta_{\boldsymbol{\rho}}^{\rm c}$ in what follows. We analyse $q_{\boldsymbol{\rho}}^{\textrm{c}}$ more thoroughly in Section 5.1, mirroring most of the discrete case experiments. In Section 5.2 we study how the results are affected when we impose the arc length parametrisation and enforce the boundary conditions to be exactly satisfied by the network $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$.